# Long-term Causal Inference via Modeling Sequential Latent Confounding

### v1

@@ 3.1 Notations, Problem Definition, Assumptions
[…]
@@
$\tau(\mathbf{x})=$ $\mathbb{E}[Y(1)-Y(0)|\mathbf{X}=\mathbf{x}]$ (4)
$=$ $\mu_{Y}^{O}(1,\mathbf{x})-\mu_{Y}^{O}(0,\mathbf{x})+\omega(\mathbf{x}),$
@@
Proof can be found in Appendix A. A similar identification result in terms of long-term average causal effects has been shown by Ghassami et al. (2022). Here we provide the identification result of HLCE $\tau(\mathbf{x})$ in Theorem 3.7.
Remark 3.8.
We illustrate Assumption 3.6 in Figure 2(a). Assumption 3.6 can be seen as a generalization of the parallel assumption in the difference-in-differences (DiD) framework (Ashenfelter & Card, 1984; Angrist & Pischke, 2009), Note that Assumption 3.6 essentially implies that the short-term confounding bias is the same as the long-term one. In this way, $\omega(\mathbf{x})$ can be interpreted as the long-term confounding bias, resulting in the identification result in Eq. (4).
@@ 4 Long-term Identification under Functional CAECB Assumption
The difference in conditional expected values of short-term potential outcomes across treated and control groups between two timesteps follows the learnable function forms $f:{\mathbb{R}}\rightarrow{\mathbb{R}}$. Formally, afor-$\forall a$, we have
@@
where $b_{t}(a,\mathbf{X})={\mathbb{E}}[S_{t}(a)|\mathbf{X},A=0,G=O]-{\mathbb{E}}[S_{t}(a)|\mathbf{X},A=1,G=O]$.
Remark 4.2.
We illustrate our proposed FCAECB assumption in Figure 2(b). We relax the CAECB assumption (Ghassami et al., 2022) to allow for the temporal short-term outcomes, instead of restricting that the short-term outcome should be the same scale as the long-term outcome. Additionally, the existing CAECB assumption can be seen as our special case when $f(\mathbf{X})$ in the FCEACB assumption satisfies $f(\mathbf{X})=1$.
@@
The proof is given in Appendix B.
Remark 4.4.
@@
$\tau(\mathbf{x})=$ $\mathbb{E}[Y(1)-Y(0)|\mathbf{X}=\mathbf{x}]$ (7)
$=$ $\mu_{Y}^{O}(1,\mathbf{x})-\mu_{Y}^{O}(0,\mathbf{x})+f^{\mu}(\mathbf{x})\omega_{T}(\mathbf{x}),$
@@
The proof is given in Appendix C. Theorem 4.5 provides the identification result of the heterogeneous long-term effects $\tau(\mathbf{X})$. The identification result consists of two parts: 1. long-term outcome differences in observational data $\mu_{S_{T}}^{O}(0,\mathbf{x})-\mu_{S_{T}}^{O}(1,\mathbf{x})$, and 2. long-term confounding bias $f^{\mu}(\mathbf{x})\omega_{T}(\mathbf{x})$. The long-term confounding bias is identified by the extrapolated result using short-term confounding bias $\omega_{T}(\mathbf{x})$ under Assumption 4.1. The identification result also illustrates a way to estimate the long-term effects $\tau(\mathbf{x})$ via modeling sequential latent confounding, which inspires our estimator as shown in Section 5.
@@ 5.1 Estimator
    Constructing final HLCE estimator as
    $\hat{\tau}(\mathbf{x})=\hat{\mu}_{Y}^{O}(1,\mathbf{x})-\hat{\mu}_{Y}^{O}(0,\mathbf{x})+\hat{f}^{\mu}(\mathbf{x})$ (9)
    ×$\times\left(\hat{\mu}_{S_{T}}^{E}(1,\mathbf{x})-\hat{\mu}_{S_{T}}^{E}(0,\mathbf{x})+\hat{\mu}_{S_{T}}^{O}(0,\mathbf{x})-\hat{\mu}_{S_{T}}^{O}(1,\mathbf{x})\right).$
Note that, in addition to the fitting/constructing steps S1-S3, we also introduce an optional S0 to select subsets of short-term outcomes $\mathbf{S}$. This step is motivated by the identification result in Eq. (7) that allows for different choices of $\mu$ and $T$. For example, suppose we can observe $6$-step short-term outcome $\mathbf{S}=[S_{1},S_{2},\dots,S_{6}]$ and the long-term outcome of interest is $Y=S_{9}$. Then we have multiple choices of $T$ and $\mu$, e.g., using all short-term outcomes $\mathbf{S}=[S_{1},S_{2},\dots,S_{6}]$ with $T=6$ and $\mu=3$, or using $\mathbf{S}=[S_{1},S_{3},S_{5}]$ with $T=3$ and $\mu=2$, or using $\mathbf{S}=[S_{1},S_{5}]$ with $T=2$ and $\mu=1$. We will discuss how to choose $T$ and $\mu$ in the next section.
@@ 5.2 Convergence Rate Analyses
In this paper, we assume the smoothness assumption of the estimated functions, where $s$-smooth functions are contained in the Hölder ball $\mathcal{H}_{d}(s)$, estimable with the minimax rate (Stone, 1980) of $n^{\frac{1}{2+d/s}}$ where $d$ is the dimension of $\mathcal{X}$. Formally, we provide the following definition.
@@
We assume that the nuisance functions and the mapping $f$ defined in Assumption 4.1 satisfy: (1) $\mu_{S_{t}}^{E}$, $\mu_{S_{t}}^{O}$, and $\mu_{Y}^{O}$ are $\alpha$-smooth, $\beta$-smooth, and $\gamma$-smooth, respectively, and all are estimable at Stone (1980)’s minimax rate of $n^{\frac{-p}{2p+d}}$ for a $p$-smooth function; (2) $f(x)$ is $\eta$-smooth.
We also assume the boundedness assumption of the nuisance functions.
@@
We now state our main theoretical results. To obtain our error bounds of the proposed estimator, we leverage the same sample splitting technique from (Kennedy, 2023), which randomly splits the datasets into two independent sets and applies them to the regressions of the first step and second step respectively. Such a technique is originally used to analyze the convergence rate of the double robust conditional average treatment effect estimation in the traditional setting (Kennedy, 2023) and later is adapted to several other methods (Curth & Van der Schaar, 2021; Frauen & Feuerriegel, ). Different from them, we use such a technique for the sequential latent confounding modeling, which is then adapted for the long-term effect estimation.
@@
Suppose the training steps S1 and S2 are train on two independent datasets of size $n$ respectively, and suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5, 4.1, and 5.2 hold, then we have
$\hat{f}(\mathbf{x})-$ $f(\mathbf{x})=O_{p}\left((\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}}\right.$ (10)
$\left.+(\frac{1}{(T-1)n})^{\frac{\alpha}{2\alpha+d}}+(\frac{1}{(T-1)n})^{\frac{\beta}{2\beta+d}}\right).$
The proof is given in Appendix D. Since $\hat{f}(\mathbf{x})$ is a time-series model, its effective sample is $(T-1)n$, thus it might achieve a faster rate if we observed a longer duration of short-term outcomes. Moreover, the rate of $\hat{f}(\mathbf{x})$ consist two part: the oracle rate $(\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}}$, and the rate of fitting nuisance functions $(\frac{1}{(T-1)n})^{\frac{\alpha}{2\alpha+d}}+(\frac{1}{(T-1)n})^{\frac{\beta}{2\beta+d}}$. If the nuisance functions $\mu_{S_{t}}^{E}$ and $\mu_{S_{t}}^{O}$ is smooth enough such that $\min\{\alpha,\beta\}\geq\eta$, then $\hat{f}(\mathbf{x})$ will attain the oracle rate.
@@
$=$ $O_{p}\left(n^{-\frac{\gamma}{2\gamma+d}}+n^{-\frac{\alpha}{2\alpha+d}}+n^{-\frac{\beta}{2\beta+d}}+\frac{\mu}{((T-1)n)^{\frac{\eta}{2\eta+d}}}\right.$
$\left.+\frac{\mu}{((T-1)n)^{\frac{\alpha}{2\alpha+d}}}+\frac{\mu}{((T-1)n)^{\frac{\beta}{2\beta+d}}}\right).$
Proof can be found in Appendix E. Theorem 5.5 follows directly from Lemma 5.4. The rate of $\hat{\tau}(\mathbf{x})$ consist of several terms: 1. $n^{-\frac{\gamma}{2\gamma+d}}$ represents the rate of $\mu_{Y}^{O}(a,\mathbf{x})$, 2. $n^{-\frac{\alpha}{2\alpha+d}}+n^{-\frac{\beta}{2\beta+d}}$ corresponds the rate of confounding bias $\omega_{T}(\mathbf{x})$, 3. and the remaining term is the rate of $\hat{f}^{\mu}(\mathbf{x})$. Theorem 5.5 suggests that the convergence rate is primarily influenced by the smoothness parameters $\alpha,\beta,\gamma,\eta$ and the temporal parameters $T,\mu$, where $T$ represents the longest observed duration of the short-term outcomes $\mathbf{S}$ and $\mu$ represents the time horizon of the long-term outcome to be estimated. This implies that achieving a more accurate effect estimation requires observing short-term outcomes for as long as possible.
@@ 6.1 Experimental Setup
In this section, we describe the data generation process for the synthetic datasets, the implementation details, and the metrics used in the experiments.
Synthetic Datasets The data generation process is partly following (Kallus et al., 2018) such that we can obtain a specific form of confounding bias. Specifically, we first generate the treatments as follows: $A|G=O\sim\text{Bernoulli}(0.6)$ and $A|G=E\sim\text{Bernoulli}(0.4)$. Then we generate the observed $\mathbf{X}$ and the unobserved $\mathbf{U}$ as follows:
$(\mathbf{X},\mathbf{U})|A,G=E\sim\mathcal{N}([\frac{2A-1}{2},0],\begin{bmatrix}1&0\\ 0&1\end{bmatrix})$ (12)
$(\mathbf{X},\mathbf{U})|A,G=O\sim\mathcal{N}([\frac{1-2A}{2},0],\begin{bmatrix}1&A-0.5\\ A-0.5&1\end{bmatrix}).$
Finally, we generate the $T$-step short-term outcomes $\mathbf{S}$ and the long-term outcome $Y=S_{\mu}$ satisfying Assumption 4.1 as follows:
$S_{t}(1)=1+1.1\mathbf{X}+\mathbf{U}+\sum_{k=0}^{t}S_{k}(1)+\epsilon_{S_{t}},$ (13)
$S_{t}(0)=\mathbf{X}+\mathbf{U}+\sum_{k=0}k^{t}S_{k}(0)+\epsilon_{S_{t}},$
where $\epsilon_{S_{t}}$ are Gaussian noises. This will result in non-equal confounding bias in different time steps $t$, i.e., $\omega_{t}(\mathbf{x})=t\times\mathbf{x}$. We perform several control experiments using this data generation process. The default values are: the observational data sample size $n_{o}=4000$, experimental data sample size $n_{e}=2000$, $T=6$, and $\mu=3$. In control experiments, we vary $\mu$ within $\{1,2,3,4,5\}$ and $T$ within $T\in\{4,5,6,7,8\}$. We also fix the ratio of $n_{O}:n_{E}=2:1$, and varying $n_{e}$ within $\{1000,2000,3000,4000,5000,10000\}$ (which corresponds to $n_{o}$ values of $\{2000,4000,6000,8000,10000,20000\}$).
Implementation We fit all nuisance functions using correctly specified regression. We denote our estimator as $\hat{\tau}$. For comparison, we evaluate our method against a correctly specified T-Learner (Künzel et al., 2019), denoted as $\hat{\tau}_{exp}$, which directly regresses unobserved $Y$ in the experimental data. However, this estimator $\hat{\tau}_{exp}$ is infeasible in practice since $Y$ is missing in the experimental data. This baseline serves as an idealized benchmark, and comparing our method against it highlights the effectiveness of our estimator across different control experiments.
@@
Metrics As for heterogeneous effect estimation, we report Precision in the Estimation of Heterogeneous Effect (PEHE) $\varepsilon_{PEHE}=\sqrt{\frac{1}{n}\Sigma_{i=1}^{n}(\tau(\mathbf{x}_{i})-\hat{\tau}(\mathbf{x}_{i}))^{2}}$. As for average long-term causal effect estimation, we report the absolute error $\varepsilon_{ATE}=|\Sigma_{i=1}^{n}\tau(\mathbf{x}_{i})-\Sigma_{i=1}^{n}\hat{\tau}(\mathbf{x}_{i}))|$. For all metrics, we report the mean values and standard errors by $100$ times running.

### v2

@@ 3.2 CAECB Assumption [14]
${\mathbb{E}}[S(a)\mid\mathbf{X},A=0,G=O]-{\mathbb{E}}[S(a)\mid\mathbf{X},A=1,G=O]$ (3)
$=$ ${\mathbb{E}}[Y(a)\mid\mathbf{X},A=0,G=O]-{\mathbb{E}}[Y(a)\mid\mathbf{X},A=1,G=O].$
@@
Suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5 and 3.6 hold, then $\tau(\mathbf{x})$ can be identified:
$\tau(\mathbf{x})=$ $\mathbb{E}[Y(1)-Y(0)\mid\mathbf{X}=\mathbf{x}]=\mu_{Y}^{O}(1,\mathbf{x})-\mu_{Y}^{O}(0,\mathbf{x})+\omega(\mathbf{x}),$ (4)
@@
Proof can be found in Appendix E. A similar identification result in terms of long-term average causal effects has been established by Ghassami et al. [14]. For completeness, we present the identification result of the HLCE $\tau(\mathbf{x})$ in Theorem 3.7.
Remark 3.8 (Interpretation on Assumption 3.6).
Assumption 3.6 is intuitively illustrated in Figure 2(a). Assumption 3.6 can be seen as a generalization of the parallel assumption in the difference-in-differences (DiD) framework [2, 3], Note that Assumption 3.6 essentially implies that the short-term confounding bias is the same as the long-term one. In this way, $\omega(\mathbf{x})$ can be interpreted as the long-term confounding bias, resulting in the identification result in Eq. (4).
@@ 4 Long-term Identification under Functional CAECB Assumption
The difference in conditional expected values of short-term potential outcomes across treated and control groups between two time steps follows the learnable function form $f:{\mathbb{R}}\rightarrow{\mathbb{R}}$. Formally, afor-$\forall a$, we have
@@
where $b_{t}(a,\mathbf{X})={\mathbb{E}}[S_{t}(a)\mid\mathbf{X},A=0,G=O]-{\mathbb{E}}[S_{t}(a)\mid\mathbf{X},A=1,G=O]$.
Remark 4.2 (Interpretation on Assumption 4.1).
We illustrate our proposed FCAECB assumption in Figure 2(b). We relax the CAECB assumption [14] to allow for the temporal short-term outcomes, instead of restricting that the short-term outcome should be the same scale as the long-term outcome. Additionally, the existing CAECB assumption can be seen as our special case when $f(\mathbf{X})$ in the FCAECB assumption satisfies $f(\mathbf{X})=1$.
@@
The proof is given in Appendix F.
Remark 4.4 (Insight of Assumption 4.7).
@@
Remark 4.5 (Testing Assumption 4.7 over Short Term Duration).
It is possible to empirically test Assumption 4.1 over the short-term duration, which provides insights into whether it can be extrapolated to long-term confounding bias. This is particularly reasonable when short- and long-term outcomes share the same practical meaning, making it more likely that the confounding bias mechanism remains stable. Specifically, suppose the short-term outcomes are observed at $3$ time steps. Then, given $S_{1},S_{2},S_{3}$, we can estimate $f(\mathbf{x})$ using $S_{1}$ and $S_{2}$, and evaluate the $R^{2}$ measure on $S_{2}$ and $S_{3}$ where $R^{2}=1-\frac{\sum_{i}^{N}(\omega_{3}(\mathbf{x}^{(i)})-\hat{f}(\mathbf{x}^{(i)})\omega_{2}(\mathbf{x}^{(i)}))^{2}}{\sum_{i}^{N}(\omega_{3}(\mathbf{x}^{(i)})-\frac{1}{N}\sum_{j}^{N}\omega_{3}(\mathbf{x}^{(j)}))^{2}}$. If $R^{2}\approx 1$, it suggests that $f$ well explains the relationships of confounding biases between adjacent time steps. Conversely, if $R^{2}\ll 1$, then the assumption may not hold, or $f$ is poorly fitted.
Example 4.6.
In many real-world scenarios, Assumption 4.1 is reasonable. In the motivating example from our introduction, AIDs patients undergo regular follow-up visits, and their short-term health indicators, such as viral loads or CD4 counts, are recorded as short-term outcomes. Since CD4 counts are critical indicators of health status, they serve as both short- and long-term outcomes. In such cases, Assumption 4.1 can be tested over short-term durations and is likely to hold.
@@
$\tau(\mathbf{x})=$ $\mathbb{E}[Y(1)-Y(0)\mid\mathbf{X}=\mathbf{x}]=\mu_{Y}^{O}(1,\mathbf{x})-\mu_{Y}^{O}(0,\mathbf{x})+f^{\mu}(\mathbf{x})\omega_{T}(\mathbf{x}),$ (7)
@@
The proof is given in Appendix G. Theorem 4.7 provides the identification result of the heterogeneous long-term effects $\tau(\mathbf{X})$. The identification result consists of two parts: 1. long-term outcome differences in observational data $\mu_{S_{T}}^{O}(0,\mathbf{x})-\mu_{S_{T}}^{O}(1,\mathbf{x})$, and 2. long-term confounding bias $f^{\mu}(\mathbf{x})\omega_{T}(\mathbf{x})$. The long-term confounding bias is identified by the extrapolated result using short-term confounding bias $\omega_{T}(\mathbf{x})$ under Assumption 4.1. The identification result also illustrates a way to estimate the long-term effects $\tau(\mathbf{x})$ via modeling sequential latent confounding, which inspires our estimator as shown in Section 5.
@@ 5.1 Estimator
    Constructing the final HLCE estimator as
    $\hat{\tau}(\mathbf{x})=\hat{\mu}_{Y}^{O}(1,\mathbf{x})-\hat{\mu}_{Y}^{O}(0,\mathbf{x})+\hat{f}^{\mu}(\mathbf{x})\left(\hat{\mu}_{S_{T}}^{E}(1,\mathbf{x})-\hat{\mu}_{S_{T}}^{E}(0,\mathbf{x})+\hat{\mu}_{S_{T}}^{O}(0,\mathbf{x})-\hat{\mu}_{S_{T}}^{O}(1,\mathbf{x})\right).$ (9)
Note that, in addition to the fitting/constructing steps S1-S3, we also introduce an optional step S0 to select appropriate subsets of short-term outcomes $\mathbf{S}$. This step is motivated by the identification result in Eq. (7) that allows for different choices of $\mu$ and $T$. For example, suppose we can observe $6$-step short-term outcome $\mathbf{S}=[S_{1},S_{2},\dots,S_{6}]$ and the long-term outcome of interest is $Y=S_{9}$. Then we have multiple choices of $T$ and $\mu$, e.g., using all short-term outcomes $\mathbf{S}=[S_{1},S_{2},\dots,S_{6}]$ with $T=6$ and $\mu=3$, or using $\mathbf{S}=[S_{1},S_{3},S_{5}]$ with $T=3$ and $\mu=2$, or using $\mathbf{S}=[S_{1},S_{5}]$ with $T=2$ and $\mu=1$. We will discuss how to choose $T$ and $\mu$ in the next section.
@@ 5.2 Convergence Rate Analyses
In this paper, we assume the smoothness assumption of the estimated functions, where $s$-smooth functions are contained in the Hölder ball $\mathcal{H}_{d}(s)$, estimable with the minimax rate [35] of $n^{\frac{1}{2+d/s}}$ where $d$ is the dimension of $\mathcal{X}$. Formally, we provide the following definition.
@@
Accordingly, we assume the smoothness assumption of the nuisance functions, as well as the boundedness assumption of the nuisance functions and their estimates as follows.
@@
We assume that the nuisance functions and the mapping $f$ defined in Assumption 4.1 satisfy: (1) $\mu_{S_{t}}^{E}$, $\mu_{S_{t}}^{O}$, and $\mu_{Y}^{O}$ are $\alpha$-smooth, $\beta$-smooth, and $\gamma$-smooth, respectively, and all are estimable at Stone [35]’s minimax rate of $n^{\frac{-p}{2p+d}}$ for a $p$-smooth function; (2) $f(x)$ is $\eta$-smooth.
@@
We now state our main theoretical results. To obtain our error bounds of the proposed estimator, we leverage the same sample splitting technique from [25], which randomly splits the datasets into two independent sets and applies them to the regressions of the first step and second step respectively. Such a technique is originally used to analyze the convergence rate of the double robust conditional average treatment effect estimation in the traditional setting [25] and later is adapted to several other methods [10, 13]. Different from them, we use such a technique for the sequential latent confounding modeling, which is then adapted for the long-term effect estimation.
@@
Suppose the training steps S1 and S2 are train on two independent datasets of size $n$ respectively, and suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5, 4.1, 5.2, and 5.3 hold, then we have
which attains the oracle rate if $\min\{\alpha,\beta\}\geq\eta$.
The proof is given in Appendix H. Since $\hat{f}(\mathbf{x})$ is a time-series model, its effective sample is $(T-1)n$, thus it might achieve a faster rate if we observed a longer duration of short-term outcomes. Moreover, the rate of $\hat{f}(\mathbf{x})$ consist two part: the oracle rate $(\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}}$, and the rate of fitting nuisance functions $(\frac{1}{(T-1)n})^{\frac{\alpha}{2\alpha+d}}+(\frac{1}{(T-1)n})^{\frac{\beta}{2\beta+d}}$. If the nuisance functions $\mu_{S_{t}}^{E}$ and $\mu_{S_{t}}^{O}$ is smooth enough such that $\min\{\alpha,\beta\}\geq\eta$, then $\hat{f}(\mathbf{x})$ will attain the oracle rate.
@@
$=$ $O_{p}\left(n^{-\frac{\gamma}{2\gamma+d}}+n^{-\frac{\alpha}{2\alpha+d}}+n^{-\frac{\beta}{2\beta+d}}+\frac{\mu}{((T-1)n)^{\frac{\eta}{2\eta+d}}}+\frac{\mu}{((T-1)n)^{\frac{\alpha}{2\alpha+d}}}+\frac{\mu}{((T-1)n)^{\frac{\beta}{2\beta+d}}}\right).$
Proof can be found in Appendix I. Theorem 5.5 follows directly from Lemma 5.4. The rate of $\hat{\tau}(\mathbf{x})$ consist of three components: (a) $n^{-\frac{\gamma}{2\gamma+d}}$ represents the rate of $\mu_{Y}^{O}(a,\mathbf{x})$; (b) $n^{-\frac{\alpha}{2\alpha+d}}+n^{-\frac{\beta}{2\beta+d}}$ corresponds the rate of confounding bias $\omega_{T}(\mathbf{x})$; (c) and the remaining term is the rate of $\hat{f}^{\mu}(\mathbf{x})$. Theorem 5.5 indicates that the overall convergence rate is largely governed by the smoothness parameters $\alpha,\beta,\gamma,\eta$ and the temporal information $T$ and $\mu$, where $T$ denotes the maximum observed duration of short-term outcomes $\mathbf{S}$, and $\mu$ is the time horizon of the long-term outcome of interest. This result aligns with intuition: both $T$ and $\mu$ affect the convergence behavior. More accurate effect estimation benefits from longer short-term outcome durations $T$, while requiring the long-term horizon $\mu$ not to be too distant.
@@ 6.1 Experimental Setup
Table 1: The results on synthetic data are presented as the mean and standard deviation. Lower values indicate better performance. The best-performing methods are bolded, with the exception of the idealized estimator $\hat{\tau}_{exp}$, which is excluded from the comparison. (See Appendix D for Full results.)
Data $\mu=1$, $T=6$ $\mu=3$, $T=6$ $\mu=5$, $T=6$
Metrics ${\varepsilon_{PEHE}}$ $\varepsilon_{ATE}$ ${\varepsilon_{PEHE}}$ $\varepsilon_{ATE}$ ${\varepsilon_{PEHE}}$ $\varepsilon_{ATE}$
$\hat{\tau}_{exp}$ (idealized) 3.0033 $\pm$ 1.8226 2.0722 $\pm$ 1.7681 12.0142 $\pm$ 7.2895 8.2901 $\pm$ 7.0709 48.0590 $\pm$ 29.1580 33.1620 $\pm$ 28.2845
T-learner (using obs. $Y$) 71.8431 $\pm$ 3.1612 6.8211 $\pm$ 2.8357 287.3693 $\pm$ 12.6434 27.2795 $\pm$ 11.3424 1149.4769 $\pm$ 50.5746 109.1205 $\pm$ 45.3685
LTEE 73.0421 $\pm$ 7.1531 10.0977 $\pm$ 6.1071 287.3264 $\pm$ 28.5202 32.4417 $\pm$ 19.5739 1168.7612 $\pm$ 122.7687 137.0583 $\pm$ 107.5567
Latent Unconfoundedness - 2.6322 $\pm$ 2.0120 - 10.5293 $\pm$ 8.0471 - 42.1179 $\pm$ 32.1874
Sequential Outcomes ($S_{\lceil T/2\rceil-1},S_{\lceil T/2\rceil},S_{\lceil T/2\rceil+1}$) - 67.5881 $\pm$ 3.0107 - 270.3607 $\pm$ 12.0428 - 1081.4361 $\pm$ 48.1759
Sequential Outcomes ($S_{1},S_{\lceil T/2\rceil},S_{T}$) - 67.5790 $\pm$ 3.0138 - 270.3237 $\pm$ 12.0553 - 1081.2884 $\pm$ 48.2252
Sequential Outcomes ($S_{1:\lceil T/2\rceil-1},S_{\lceil T/2\rceil},S_{\lceil T/2\rceil+1:T}$) - 67.5797 $\pm$ 3.0151 - 270.3266 $\pm$ 12.0607 - 1081.3001 $\pm$ 48.2472
NNPIV_LU 70.7124 $\pm$ 2.5320 6.1302 $\pm$ 2.8888 282.8595 $\pm$ 10.1320 24.5293 $\pm$ 11.5558 1131.4331 $\pm$ 40.5209 98.1118 $\pm$ 46.2263
NNPIV_SInd 70.7155 $\pm$ 2.5329 6.1306 $\pm$ 2.8884 282.8655 $\pm$ 10.1314 24.5240 $\pm$ 11.5558 1131.4602 $\pm$ 40.5252 98.0989 $\pm$ 46.2232
CAECB (using first $S$) 70.7257 $\pm$ 3.1153 6.7160 $\pm$ 2.7984 286.2518 $\pm$ 12.5973 27.1744 $\pm$ 11.3048 1148.3595 $\pm$ 50.5286 109.0154 $\pm$ 45.3308
CAECB (using middle $S$) 67.3754 $\pm$ 2.9827 6.3982 $\pm$ 2.6884 282.9014 $\pm$ 12.4626 26.8566 $\pm$ 11.1910 1145.0089 $\pm$ 50.3933 108.6976 $\pm$ 45.2162
CAECB (using last $S$) 36.1202 $\pm$ 2.0211 3.5778 $\pm$ 1.8603 251.6244 $\pm$ 11.2383 23.8972 $\pm$ 10.2153 1113.7291 $\pm$ 49.1287 105.7381 $\pm$ 44.1729
CAECB (using random $S$) 59.8772 $\pm$ 11.9929 5.7194 $\pm$ 2.6871 275.4020 $\pm$ 16.0775 26.1778 $\pm$ 10.9897 1137.5094 $\pm$ 50.1542 108.0188 $\pm$ 44.9689
Ours $\hat{\tau}$ 3.1110 $\pm$ 1.8230 2.1470 $\pm$ 1.7826 13.9945 $\pm$ 7.1682 9.296 $\pm$ 7.0820 64.3076 $\pm$ 32.9868 40.5425 $\pm$ 29.7264
@@
In this section, we describe the data generation process for the synthetic datasets, the baselines, the implementation details, and the metrics used in the experiments.
Synthetic Datasets We partially follow the data generation setup in [24] to induce a specific form of confounding bias and define nuisance functions. Details are provided in Appendix D. The confounding bias at time $t$ is designed as $\omega_{t}(\mathbf{x})=-2^{t}\mathbf{x}$, aligning with our assumptions. By default, we set the observational sample size $n_{o}=4000$, experimental sample size $n_{e}=2000$, and $T=6$, $\mu=3$. In variations, we change $\mu\in{1,2,3,4,5}$, $T\in{4,5,6,7,8}$, and fix the ratio $n_{o}:n_{e}=2:1$ while varying $n_{e}\in\{1000,2000,3000,4000,5000,10000\}$ (implying $n_{o}\in\{2000,4000,6000,8000,10000,2000\}$).
Baselines and Implementation We compare our method against several baselines for long-term causal effect estimation, including LTEE [9], Latent Unconfoundedness [5], Sequential Outcomes [20], NNPIV [28], and CAECB [14]. For the Sequential Outcomes method, which requires three groups of short-term outcomes ($S_{1},S_{2},S_{3}$ in their paper), we test three grouping strategies: (1) $S_{\lceil T/2\rceil-1},S_{\lceil T/2\rceil},S_{\lceil T/2\rceil+1}$, (2) $S_{1},S_{\lceil T/2\rceil},S_{T}$, and (3) $S_{1:\lceil T/2\rceil-1},S_{\lceil T/2\rceil},S_{\lceil T/2\rceil+1:T}$. NNPIV is applied under both the latent unconfoundedness assumption (NNPIV_LU) [16] and the surrogacy assumption (NNPIV_SInd) [4]. For CAECB, which uses a single short-term outcome, we implement four variants using the first, middle, last, and a randomly selected outcome. We also include the T-learner [26] trained on observational $Y$ and an idealized version trained on experimental $Y$ (infeasible in practice, as $Y$ is unobserved in experimental data), serving as an idealized benchmark denoted as $\hat{\tau}_{exp}$. All nuisance functions are estimated using correctly specified regression. Our estimator is denoted as $\hat{\tau}$. Additional details are provided in Appendix D.
@@
Metrics As for heterogeneous effect estimation, we report Precision in the Estimation of Heterogeneous Effect (PEHE) $\varepsilon_{PEHE}=\sqrt{\frac{1}{n}\Sigma_{i=1}^{n}(\tau(\mathbf{x}_{i})-\hat{\tau}(\mathbf{x}_{i}))^{2}}$. As for average long-term causal effect estimation, we report the absolute error $\varepsilon_{ATE}=|\Sigma_{i=1}^{n}\tau(\mathbf{x}_{i})-\Sigma_{i=1}^{n}\hat{\tau}(\mathbf{x}_{i}))|$. For all metrics, we report the mean values and standard deviation (in tables) or standard errors (in figures) by $50$ running.